# AttentionStore: Cost-effective Attention Reuse across Multi-turn Conversations in Large Language Model Serving

### v2

@@ 4.3 Ablation Studies
4.3.1 Recomputation v.s. AttentionStore
We investigate the prefilling performance of different methods under varying historic and new token ratios. Different methods prefill the same 1K tokens under the batch size of 16 on an A100 GPU for LLaMA-13B. RE computes the KV cache for all tokens, while AS loads the KV cache of historical tokens from the KV caching system and partially prefills the new input tokens. For example, the setting 600/400 means AS loads the KV cache of 600 tokens and computes the KV cache for 400 tokens. Overall, AS outperforms RE in all tested settings, as shown in Figure 18. This advantage becomes more pronounced as the percentage of newly input tokens decreases (from 500 to 100), as depicted by the middle bar of each bar group. Although the KV cache loading time for AS gradually increases with the percentage of historical tokens (from 500 to 900), the layer-wise pre-loading scheme effectively eliminates this loading time, as demonstrated by the third bar of each bar group. Note that when the KV cache loading time exceeds the prefilling time (e.g., the second bar of setting 900/100), AS can conceal the KV cache loading time by enabling a read buffer. The impact of the read buffer is evaluated in the next subsection.
@@ 4.3.2 Overlapped KV cache Access
Figure 19: AS with no pre-loading v.s. AS pre-loading with various buffer sizes.
@@
Layer-wise KV cache pre-loading. In the experiments, we set the length of historical tokens to 1K and the length of newly input tokens to 100 for investigating the effectiveness of the lay-wise pre-loading scheme. The first bar in Figure 20, i.e., NO-PL, shows the time of prefilling without the pre-loading scheme that includes two parts: the KV cache loading time and the computation time of the newly input tokens. The following bars in Figure 20 show the prefilling time when the layer-wise pre-loading scheme has different sizes of read buffers. For clarity, we use the number of layers to represent the buffer size, e.g., PL-B0 indicates no read buffer and PF-B5 indicates a read buffer size of 5 layers of KV cache. We observe although there is no read buffer, i.e., PL-B0, the pre-loading scheme reduces the prefilling time by 35% compared to NO-PL. PF-B15 perfectly overlaps the KV cache loading time and reduces the prefilling time by 61% compared to NO-PL.
Asynchronous KV cache saving. In the experiments, we set different prompt lengths ranging from 1K to 1.6K and the number of decoding steps to 20 for investigating the effectiveness of the asynchronous saving scheme. As shown in Figure 20, we observe that the saving time increases as the prompt length grows, since the size of the KV cache to be saved increases. To mitigate the saving overhead, AttentionStore employs the asynchronous saving scheme that allows the KV cache saving to overlap with the execution of the inference, reducing the overall execution time by 13% to 15%.
@@ 4.3.3 Scheduler-aware Fetching and Eviction
We investigate the effectiveness of the scheduler-aware fetching and eviction in AttentionStore upon improving the cache hit rate. We compare the overall cache hit rates, DRAM hit rates, and disk hit rates of AS and existing eviction policies (including LRU and FIFO) across various storage configurations.
As shown in Figure 21(a), for the configuration of 128G/2T that indicates 128GB DRAM and 2TB SSD, AS outperforms LRU and FIFO in the overall cache hit rate by 27% and 31%, respectively. With the increased SSD capacity (128G/10T), AS achieves a remarkable hit rate of 86%, surpassing LRU (58%) and FIFO (48%). AS achieves high overall hit rates because AS is aware of the future KV cache access information to avoid evicting the KV caches that will be used in the future. The higher hit rates are translated to the reduced GPU time as shown in Figure 21(b), where AS achieves speedup up to 2.7$\times$. Analyzing the breakdown of hit rate, for the configuration of 128G/2T, LRU and FIFO only achieve 0.5% and 0.5% (too tiny to display) DRAM hit rates, with the remaining 12.4% and 9.0% being disk hit rates. Even with the overall hit rate increasing to 58% and 48% respectively for the larger capacity of 128G/10T, LRU and FIFO still exhibit limited DRAM hit rates of approximately 0.6% and 0.5% respectively. This is because LRU and FIFO lack awareness of future KV cache information and cannot pre-fetch KV caches from disks to host memory, thereby limiting their ability to improve DRAM hit rates. In contrast, AS achieves a cache hit rate of up to 86%, with over 99.6% of the hits occurring in DRAM due to its scheduler-aware policy.
@@ 4.3.4 Performance of Decoupled KV Cache Truncation
When the context window exceeds its limit, AttentionStore truncates the KV cache directly, thus avoiding the need for re-computation and reducing overhead. We evaluate the effectiveness of the way AS used to manage the context overflow. Specifically, we compare a baseline approach overflow (OF) that embeds positional encoding within the KV caches, leading to the invalidation of KV caches in the KV caching system. OF relegates context overflow to be managed by recomputation. This experiment uses 128GB DRAM and 10TB SSD. As evident from Figure 22(a), comparing OF with AS, the hit rates decrease by 17.6%, 41.5%, 18.1%, 18.4% for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively. This decline is attributed to the fact that if applying OF, every instance of context overflow necessitates context truncation, thereby invalidating the KV caches in the KV caching system. This decrease in hit rate subsequently translates to the longer GPU time as shown in Figure 22(b).
AttentionStore ensures the validity of the saved KV caches in the system when context overflow happens and promises a higher hit rate and reduced GPU time. OF of LLaMA-65B experiences a low hit rate due to its limited 2K context window. After serving a conversation of the first turn, the session easily reaches the context window limit, consequently making the associated KV cache invalidate. Subsequently, the following conversations in the same session face KV cache miss.
4.3.5 The Cache Capacity Requirement
(a) Impact on hit rate.
(b) Impact on throughput.
In this subsection, we investigate how much cache capacity AttentionStore needs to achieve a remarkable cache hit rate. The cache capacity required is related to the maximum number of distinct conversation sessions served by an LLM serving system per unit time (denoted as $DSpUT$). The larger the $DSpUT$ value is, the more distinct sessions the system handles per unit time, and the more KV cache storage space is required. Moreover, due to the limitation of the maximum context window, the maximum KV cache capacity required by one conservation session is fixed, i.e., equal to the length of the maximum context window multiplied by the KV size of each token, which is denoted as $CCpS$. Thus the required maximum cache capacity per unit time is $CCpUT=DSSpUT*CCpS$. In AttentionStore, the KV cache of each session has a TTL (time to live) that indicates its maximum saving time since the last access. The TTL is set as the unit time mentioned above. By configuring the cache capacity of AttentionStore as $CCpUT$, we can achieve a cache hit rate of 100% if not considering the newly arrived conversations. Nevertheless, to achieve a high cache hit rate in real scenarios, we do not need to configure such a large capacity since the hotness of cached items is different.
4.3.6 Impact of Caching Storage Mediums
(a) Impact of caching storage mediums on hit rates.
(b) Impact of caching storage mediums on GPU time.
Figure 24: Performance under various caching configurations.
Some existing works [67, 19, 30] employ only the HBM space for caching the KV caches of multi-turn conversations. We here compare the performance of mechanism caching KVs on HBMs with that of AttentionStore caching KVs on DRAM and SSDs. In the experiments, we configure the size of the HBM cache as 10GB, the size of DRAM as 128GB, and the size of SSDs as 10TB. Figure 24(a) shows cache hit rates and inference performance of different mechanisms. The hit rate of the HBM-only caching method is nearly 0% for all models due to the limited capacity of HBM. Using HBM with DRAM improves the cache hit rate to 3.4%, 1.7%, 7.7%, and 19.1% for models LLaMA-13B, LLaMA-65B, LLaMA-70B and Falcon-40B, respectively. In contrast, by further extending the cache capacity with SSDs, AttentionStore improves the cache hit rate to 86%, 71%, 89%, and 90% for models LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively. With higher hit rates, AttentionStore significantly improves the inference performance compared to the HBM-only/HBM+DRAM policies as shown in Figure 24(b).
@@ 4.3.7 Accuracy of Decoupled Positional Encoding
To maintain the validity of the stored KV caches, AttentionStore decouples positional encoding from the KV caches and embeds the new positional encodings when reusing the stored KV caches as presented in Section 3.4. We evaluate the impact of the different schemes including AS, token truncation (TT), and naive KV cache truncation (NKVT) on the perplexity (PPL) and the accuracy of LLMs leveraging widely used benchmarks. In situations where the number of historical tokens exceeds the context window limit, TT removes the historical tokens and recomputes the KV caches for the remaining tokens, and NKVT directly discards the KV caches associated with the positional encoding and utilizes the truncated KV caches instead.
PPL. PPL is a metric used to evaluate the quality of a model in generating tokens [58, 14]. Lower PPL values indicate that the model is better at predicting the text and demonstrates a greater understanding of the language. Table 1 shows the PPL comparison of LLaMA-7B and LLaMA-13B in AS, TT, and NKVT settings using datasets WikiText-2 [15], C4 [39], and PTB [27]. TT consistently achieves a low PPL by recomputing KV caches after context window overflow. AS also maintains a low PPL, comparable to TT (with a difference of $<0.02$), by incorporating new positional encodings into the KV caches after truncation. In contrast, NKVT exhibits a high PPL ($>10^{3}$) due to the coupling of positional encoding within its KV caches. Directly truncating the KV caches would scramble the coupled positional information, resulting in the models’ failure to maintain a low PPL.
Accuracy. To analyze the accuracy of the models in answering questions after truncation, we conduct experiments using the MMLU [18], LongEval [23, 58], and PIQA [7] benchmarks. Specifically, we first input a long text to simulate the overflow of historical inputs to trigger the truncation operation, and then append the questions from the benchmarks as new inputs afterward. As shown in Table 2, both AS and TT provide high comparable accuracy. TT achieves high accuracy by paying the recomputation cost for context window overflow, while our AS avoids this cost and still maintains high accuracy. In contrast, the NKVT has a much lower accuracy than AS and TT because the coupled positional encoding after KV cache truncation is miscoded, which results in more disruption to new inputs.
@@
Dataset Model AS TT NKVT
@@
Benchmark Model AS TT NKVT
@@
4.3.8 Performance for Long Sequence Inference
(a) Impact on TTFT.
(b) Impact on prefill throughput.
Figure 25: Prefilling performance of long sequence inference.
Modern LLMs continue to incorporate longer context windows to accommodate a greater amount of information, empowering long sequence inference applications (e.g., document understanding [60] and code understanding [29]). We assess the efficacy of AttentionStore with models designed for long sequence inference in these applications. Specifically, we deploy the Mistral-7B model [20] with a maximum 32K context window on one A100 GPU with 80GB HBM, employing a GQA factor of 8 [2, 20]. We evaluate a documentation analysis application as an example. In this application, users submit a series of analysis tasks for the same document, forming a multi-turn conversation session. The size of the document varies from 4K to 28K. Each session consists of 6 analysis tasks, with each task requiring an input of 256 tokens and producing an output of 64 tokens. In the experiments, we use a batch size of 1 and evaluate the performance of the second and subsequent turns.
TTFT. Figure 25(a) shows the average TTFT for different sequence lengths. The TTFT of RE gradually increases to 2.5s when the sequence length reaches 28K. In contrast, AS has only about 0.12s of TTFT, resulting in a 95% reduction in TTFT compared to RE.
Prefilling throughput. Figure 25(b) illustrates the measured prefilling throughput. AS significantly improves the prefilling throughput, achieving a speedup up to $\times2$ compared to RE. The prefilling throughput of RE cannot improve as the sequence length grows due to being bounded by the computational capability. AS observes a continuous increase in the prefilling throughput as the sequence length grows by efficiently reusing the historical KV cache.
GPU time. Figure 26(a) shows the average GPU time to complete each analysis task. AS demonstrates consistent GPU time savings compared to RE, regardless of the sequence length. When the sequence length increases, RE requires more prefilling time to recompute the KV caches, which accounts for a substantial portion of the total GPU time, i.e., 41%, for a sequence length of 28K. In contrast, AS efficiently reduces prefilling costs by reusing KV caches, resulting in only 1.2% of the GPU time being allocated to prefilling.
Output throughput. The overall output throughput is calculated by the number of generated tokens divided by the total processing time of a task, as shown in Figure 26(b). As the sequence length increases, both RE and AS experience a decrease in throughput. This is attributed to the increased computational demands for computing attention of the lengthy sequence, resulting in a longer decoding time for each token. Notably, AS consistently surpasses RE in all scenarios, demonstrating an improvement in output throughput of up to 67%. These improvements in output throughput primarily stem from the elimination of KV cache recomputation.
(a) Impact on GPU time.
(b) Impact on output throughput.
Figure 26: Overall performance of long sequence inference.
@@ 5 Related Work
KV Cache Management. Within a single-turn conversation, the KV cache is widely used for improving the performance of the decoding phase [52, 61, 67, 10, 45, 54]. To reduce the storage overhead of the KV cache on HBMs, existing work employs quantization and compression techniques on KV caches [58, 13, 65, 25]. To reduce the memory waste incurred by fragmentation, vLLM [22] takes inspiration from virtual memory to allow the KV cache to use fine-granularity non-continuous memory. These techniques are orthogonal to AttentionStore, which focuses on multi-turn conversations.
LMDeploy [19] is an LLM inference framework that caches the KV caches of multi-turn conversations on HBMs. RadixAttention [67], ChunkAttention [60], and Pensieve [62] are inference techniques that were developed concurrently with AttentionStore. RadixAttention and ChunkAttention optimize the inference tasks that share prompt prefixes. Tasks with the same prompt prefixes share the same KV caches to reduce the KV computation. Pensieve utilizes both GPU and CPU memory to store KV caches for multi-turn conversations. Different from all these works, AttentionStore exploits slower but larger storage hierarchies to save the KV caches to achieve high cache hit rates as presented in Section 4.3.6, and focuses on designing systemic techniques to address the challenges of offloading to slower mediums.
Inference Parameter Offloading. FlexGen [43] offloads both model weights and KV cache to DRAM and disks to support offline inference of LLMs. DeepSpeed Inference [4, 41, 40] offloads model weights to the DRAM and disks and fetches them on demand. Lina [24] offloads infrequently used expert weights of LLMs to the host memory to improve the memory efficiency. PowerInfer [44] and LLM in a flash[3] utilize sparsity [26, 28] in FFN computation to offload most of the inactive weights to the host memory or disks to reduce both memory usage and the computation. FastServe [56] schedules the KV caches to the host memory for optimizing the job completion time. In contrast, AttentionStore exploits KV caching offloading to reduce the recomputation overhead in multi-turn conversations.

### v3

@@ 4.3 Ablation Studies
4.3.1 Recomputation v.s. CachedAttention
We investigate the prefilling performance of different methods under varying historic and new token ratios. Different methods prefill the same 1K tokens under the batch size of 16 on an A100 GPU for LLaMA-13B. RE computes the KV cache for all tokens, while CA loads the KV cache of historical tokens from AttentionStore and partially prefills the new input tokens. For example, the setting 600/400 means CA loads the KV cache of 600 tokens and computes the KV cache for 400 tokens. As shown in Figure 18, CA consistently outperforms RE in all tested settings. This advantage becomes more pronounced as the percentage of newly input tokens decreases (from 500 to 100), as depicted by the middle bar of each bar group. Although the KV cache loading time for CA gradually increases with the percentage of historical tokens (from 500 to 900), the layer-wise pre-loading scheme effectively eliminates this loading time, as demonstrated by the third bar of each bar group. Note that when the KV cache loading time exceeds the prefilling time (e.g., the second bar of setting 900/100), CA can conceal the KV cache loading time by enabling a read buffer. The impact of the read buffer is evaluated in the next subsection.
@@ 4.3.2 Overlapped KV cache Access
Figure 19: CA with no pre-loading v.s. CA pre-loading with various buffer sizes.
@@
Layer-wise KV cache pre-loading. We set the length of historical tokens to 1K and the length of newly input tokens to 100 for investigating the effectiveness of the lay-wise pre-loading scheme. The first bar in Figure 20, i.e., NO-PL, shows the time of prefilling without the pre-loading scheme that includes two parts: the KV cache loading time and the computation time of the newly input tokens. The following bars in Figure 20 show the prefilling time when the layer-wise pre-loading scheme has different sizes of read buffers. For clarity, we use the number of layers to represent the buffer size, e.g., PL-B0 indicates no read buffer and PF-B5 indicates a read buffer size of 5 layers of KV cache. We observe although there is no read buffer, i.e., PL-B0, the pre-loading scheme reduces the prefilling time by 35% compared to NO-PL. PF-B15 perfectly overlaps the KV cache loading time and reduces the prefilling time by 61% compared to NO-PL.
Asynchronous KV cache saving. We set different prompt lengths ranging from 1K to 1.6K and the number of decoding steps to 20 for investigating the effectiveness of the asynchronous saving scheme. As Figure 20 shows, the saving time increases as the prompt length grows, as the size of the KV cache to be saved increases. To reduce the saving overhead, CachedAttention employs the asynchronous saving scheme allowing KV cache saving to overlap with inference execution, reducing the overall execution time by 13% to 15%.
@@ 4.3.3 Scheduler-aware Fetching and Eviction
We evaluate the effectiveness of the scheduler-aware fetching and eviction in CachedAttention upon improving the cache hit rate. We compare overall cache hit rates, DRAM hit rates, and disk hit rates of CA and existing eviction policies (including LRU and FIFO) across various storage configurations.
As shown in Figure 21(a), for the configuration of 128G/2T that indicates 128GB DRAM and 2TB SSD, CA outperforms LRU and FIFO in the overall cache hit rate by 27% and 31%, respectively. With the increased SSD capacity (128G/10T), CA achieves a remarkable hit rate of 86%, surpassing LRU (58%) and FIFO (48%). CA achieves high overall hit rates because CA is aware of the future KV cache access information to avoid evicting the KV caches that will be used in the future. The higher hit rates are translated to the reduced GPU time as shown in Figure 21(b), where CA achieves speedup up to 2.7$\times$. Analyzing the breakdown of hit rate, for the configuration of 128G/2T, LRU and FIFO only achieve 0.5% and 0.5% (too tiny to display) DRAM hit rates, with the remaining 12.4% and 9.0% being disk hit rates. Even with the overall hit rate increasing to 58% and 48% respectively for the larger capacity of 128G/10T, LRU and FIFO still exhibit limited DRAM hit rates of approximately 0.6% and 0.5% respectively. This is because LRU and FIFO lack awareness of future KV cache information and cannot pre-fetch KV caches from disks to host memory, thereby limiting their ability to improve DRAM hit rates. In contrast, CA achieves a cache hit rate of up to 86%, with over 99.6% of the hits occurring in DRAM due to its scheduler-aware policy.
@@ 4.3.4 Performance of Decoupled KV Cache Truncation
When the context window exceeds its limit, CachedAttention truncates the KV cache directly, thus avoiding the need for re-computation and reducing overhead. We evaluate the effectiveness of the way to manage the context overflow in CA. Specifically, we compare a baseline approach overflow (OF) that embeds positional encoding within the KV caches, leading to the invalidation of KV caches in AttentionStore. OF relegates context overflow to be managed by recomputation. This experiment uses 128GB DRAM and 10TB SSD. As evident from Figure 22(a), comparing OF with CA, the hit rates decrease by 17.6%, 41.5%, 18.1%, 18.4% for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively. This decline is attributed to the fact that if applying OF, every instance of context overflow necessitates context truncation, thereby invalidating the KV caches in AttentionStore. This decrease in hit rate subsequently translates to the longer GPU time as shown in Figure 22(b).
CachedAttention ensures the validity of the saved KV caches in the system when context overflow happens and promises a higher hit rate and reduced GPU time. OF of LLaMA-65B experiences a low hit rate due to its limited 2K context window. After serving a conversation of the first turn, the session easily reaches the context window limit, consequently making the associated KV cache invalid. Subsequently, the following conversations in the same session face KV cache miss.
@@ 4.3.5 Accuracy of Decoupled Positional Encoding
To maintain the validity of the stored KV caches, CachedAttention decouples positional encoding from the KV caches and embeds the new positional encodings when reusing the stored KV caches as presented in Section 3.4. We evaluate the impact of the different schemes including CA, token truncation (TT), and naive KV cache truncation (NKVT) on the perplexity (PPL) and the accuracy of LLMs leveraging widely used benchmarks. In situations where the number of historical tokens exceeds the context window limit, TT removes the historical tokens and recomputes the KV caches for the remaining tokens, and NKVT directly discards the KV caches associated with the positional encoding and utilizes the truncated KV caches instead.
PPL. PPL is a metric used to evaluate the quality of a model in generating tokens [54, 13]. Lower PPL values indicate that the model is better at predicting the text and demonstrates a greater understanding of the language. Table 1 shows the PPL comparison of LLaMA-7B and LLaMA-13B in CA, TT, and NKVT settings using datasets WikiText-2 [14], C4 [35], and PTB [26]. TT consistently achieves a low PPL by recomputing KV caches after context window overflow. CA also maintains a low PPL, comparable to TT (with a difference of $<0.02$), by incorporating new positional encodings into the KV caches after truncation. In contrast, NKVT exhibits a high PPL ($>10^{3}$) due to the coupling of positional encoding within its KV caches. Directly truncating the KV caches would scramble the coupled positional information, resulting in the models’ failure to maintain a low PPL.
Accuracy. To analyze the accuracy of the models in answering questions after truncation, we conduct experiments using the MMLU [17], LongEval [22, 54], and PIQA [6] benchmarks. Specifically, we first input a long text to simulate the overflow of historical inputs to trigger the truncation operation, and then append the questions from the benchmarks as new inputs afterward. As shown in Table 2, both CA and TT provide high comparable accuracy. TT achieves high accuracy by paying the recomputation cost for context window overflow, while our CA avoids this cost and still maintains high accuracy. In contrast, the NKVT has a much lower accuracy than CA and TT because the coupled positional encoding after KV cache truncation is miscoded, which results in more disruption to new inputs.
@@
Dataset Model CA TT NKVT
@@
Benchmark Model CA TT NKVT
@@
4.3.6 The Cache Capacity Requirement
(a) Impact on hit rate.
(b) Impact on throughput.
[…]
4.3.7 Impact of Caching Storage Mediums
(a) Impact of caching storage mediums on hit rates.
(b) Impact of caching storage mediums on GPU time.
Figure 24: Performance under various caching configurations.
Some existing works [63, 18, 28] employ only the HBM space for caching the KV caches of multi-turn conversations. We here compare the performance of mechanism caching KVs on HBMs with that of CachedAttention caching KVs on DRAM and SSDs. In the experiments, we configure the size of the HBM cache as 10GB, the size of DRAM as 128GB, and the size of SSDs as 10TB. Figure 24(a) shows cache hit rates and inference performance of different mechanisms. The hit rate of the HBM-only caching method is nearly 0% for all models due to the limited capacity of HBM. Using HBM with DRAM improves the cache hit rate to 3.4%, 1.7%, 7.7%, and 19.1% for models LLaMA-13B, LLaMA-65B, LLaMA-70B and Falcon-40B, respectively. In contrast, by further extending the cache capacity with SSDs, CachedAttention improves the cache hit rate to 86%, 71%, 89%, and 90% for models LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively. With higher hit rates, CachedAttention significantly improves the inference performance compared to the HBM-only/HBM+DRAM policies as shown in Figure 24(b).
(a) Hit rates.
(b) TTFT.
(c) Prefill TP.
(d) GPU time.
Figure 25: Impact of session arrival rates.
4.3.8 Impact of Session Arrival Rates
We evaluate how the session arrival rates affect the performance of the CachedAttention in terms of hit rates, TTFT, prefilling throughput, and GPU time. In the experiments, we configure the session arrival rates varying from 0.5/s to 2.0/s. We test LLaMA-13B with 128GB/10TB storage following warmup as before. As shown in Figure 25(a), the cache hit rate has a slight drop down, from 82% to 77%, as the session arrival rates increase from 0.5 to 2. If the session arrival rates are low, a specific time window will contain fewer sessions, indicating a smaller DSpUT discussed in Section 4.3.6. The DSpUT at low session arrival rates is less than that at high session arrival rates. For equivalent hit rates, a lower DSpUT necessitates less cache storage space. In other words, given the same cache storage, a lower DSpUT results in higher hit rates, as the results observed in Figure 25(a). Consequently, due to the KV cache being less hit, more sessions need to recompute the KV cache, leading to an increase in the TTFT from 0.122s to 0.154s, as shown in Figure 25(b). The prefilling throughput shown in Figure 25(c) also experiences a decline, dropping from 858K/s to 681K/s, as an outcome of the slightly decreased KV cache hits. The GPU time shown in Figure 25(d) escalates from 6.25H to 7.01H. Overall, the impact of varying session arrival rates has a minimal impact on the effectiveness. Note that even with an increase in session arrival rates, the KV cache hit rates remain significant, implying that CachedAttention can still effectively leverage the KV cache, consistently offering substantial performance improvement over recomputation.
@@ 5 Related Work
KV Cache Management. Within a single-turn conversation, the KV cache is widely used for improving the performance of the decoding phase [48, 57, 63, 9, 41, 50]. To reduce the storage overhead of the KV cache on HBMs, existing work employs quantization and compression techniques on KV caches [54, 12, 61, 24]. To reduce the memory waste incurred by fragmentation, vLLM [21] takes inspiration from virtual memory to allow the KV cache to use fine-granularity non-continuous memory. These techniques are orthogonal to CachedAttention, which focuses on multi-turn conversations.
LMDeploy [18] is an LLM inference framework that caches the KV caches of multi-turn conversations on HBMs. RadixAttention [63], ChunkAttention [56], and Pensieve [58] are inference techniques that were developed concurrently with CachedAttention. RadixAttention and ChunkAttention optimize the inference tasks that share prompt prefixes. Tasks with the same prompt prefixes share the same KV caches to reduce the KV computation. Pensieve utilizes both GPU and CPU memory to store KV caches for multi-turn conversations. Different from all these works, CachedAttention exploits slower but larger storage hierarchies to save the KV caches to achieve high cache hit rates as presented in Section 4.3.7, and focuses on designing systemic techniques to address the challenges of offloading to slower mediums.
Inference Parameter Offloading. FlexGen [39] offloads both model weights and KV cache to DRAM and disks to support offline inference of LLMs. DeepSpeed Inference [3, 37, 36] offloads model weights to the DRAM and disks and fetches them on demand. Lina [23] offloads infrequently used expert weights of LLMs to the host memory to improve memory efficiency. PowerInfer [40] and LLM in a flash[2] utilize sparsity [25, 27] in FFN computation to offload most of the inactive weights to the host memory or disks to reduce both memory usage and the computation. FastServe [52] schedules the KV caches to the host memory for optimizing the job completion time. In contrast, CachedAttention exploits KV cache offloading to reduce the recomputation overhead in multi-turn conversations.